A PULLEY FIXED TO THE CEILING CARRIES A STRING WITH BLOCKS OF MASS m AND 3m. THE MASSES OF PULLEY AND STRING ARE NEGLIGIBLE.  WHAT IS THE ACCELERATION OF THE CENTRE OF MASS? 
   
OPTIONS:
1. -g/2
2.-g/4
3. g

Grade:11

2 Answers

Arun
4 years ago
Let T be the tension in the string carrying the masses m and 3m.
Let a be the acceleration, then
T -mg = ma
3mg – T = 3ma
2mg = 4ma
a = g/2
Khimraj
4 years ago
a = (3m-m)g/4m = g/2
ACCELERATION OF THE CENTRE OF MASS = (3m*g/2 – m*g/2)/4m = g/4
....................................